# Safe Text-to-Image Generation: Simply Sanitize the Prompt Embedding

### v1

@@ 5.2 Effectiveness and Robustness Evaluation
As shown in Table III, we first report the total number of inappropriate content generated by the base model (SD-v1.4 without any safeguards) on each dataset. For example, the original SD-v1.4 generates 91 instances of hateful content using the I2P dataset. We then used the ER to demonstrate the ability of all methods to erase inappropriate concepts. The results show that ES achieves SOTA level of effectiveness, with the best erasure performance for all misclassifications in the I2P dataset. Notably, since SC and SafeGen are only effective for explicit content, we evaluate their performance only in that category. For erasure of sexual content in the I2P dataset, ES achieves the highest ER of 93.03%, while the lowest among the baselines is CA at 16.02% and the highest is ESD at 72.70%. In terms of the total erasure effect across all inappropriate categories, ES achieves the highest average ER of 81.91%, followed by NP with an average ER of 56.53%. This indicates that ES surpasses the current best baselines in effectiveness by approximately 40%.
@@
Additionally, ES also achieves SOTA robustness, showing the best average defense performance against three types of adversarial attacks. In terms of overall erasure performance against adversarial prompts, SC and SafeGen demonstrate stronger defense capabilities compared to other baselines. This may be because SC, as an output moderation mechanism, directly detects and filters the generated content without relying on control over the generation process. For SafeGen, its text-agnostic design makes it less affected by synonym transformations of target concepts. In contrast, ES achieves the highest average ER of 95.04%, indicating excellent synonym generalization ability for target concepts and effective resistance to inappropriate handcrafted or adversarial prompts.
Figure 4 further demonstrates the defense effectiveness of the baseline methods and ES on the RAB adversarial prompt dataset. We use NudeNet[23] to detect exposed content in the generated images, which can identify various categories of exposure, such as female breasts (F_Breast) and male genitalia (M_Genitalia). It is noteworthy that the RAB attack generates more explicit content compared to I2P by adding extra pornographic semantics to the target prompt (the number of exposed content increases from 674 to 3596). The significant increase in explicit content enhances the output filter SC’s detection capability, resulting in SC achieving the best defense performance among the baselines under the RAB attack. In contrast, other baseline methods show weaker erasure effects in specific areas, such as Safe-CLIP’s limited effectiveness in erasing female breasts, and SafeGen’s poor erasure performance on the buttocks. However, ES effectively suppresses the generation of all exposed body parts, with an ER of 98.92% for total exposed areas, achieving SOTA performance. In addition, the visual comparison between ES and the baseline methods provided in Figure 5 further validates the effectiveness of ES.
@@
Figure 5: Visual comparison of generated content between ES and six safe generation baselines. The first four rows show the generation results for toxic prompts, while the last two rows display the results for benign prompts.
5.3 Specificity and Fidelity Evaluation
We use CLIP scores and FID scores to evaluate the specificity and fidelity of ES, comparing it with the base model SD-v1.4 without any safeguards and six safe generation baselines. In the experiments, we use the COCO-2017[18] validation set as the real image set to support the calculation of FID scores. This validation set consists of 5,000 clean image-text pairs, where each image is annotated with a corresponding description. We use these annotations as generation prompts to produce 5,000 clean images and fix the same random seed for each prompt to ensure fair comparisons, making the generated image set consistent in content and quantity with the real image set for comprehensive quality evaluation.
@@ 5.3 Specificity and Fidelity Evaluation
Table IV reports the quantitative results of each method in terms of specificity and fidelity. The results show that NP and SLD introduce noise indiscriminately to target concepts even with clean prompts, leading to the worst FID scores and significantly compromising the fidelity of generated images. Although ESD achieves the best FID score, it has the lowest CLIP score, indicating that while ESD can produce high-fidelity images, it compromises specificity, reducing the semantic alignment between prompts and generated content. This observation is confirmed by the visual results in Figure 5. CA performs well in both specificity and fidelity, likely at the cost of effectiveness and robustness. SafeGen, in comparison, suppresses only concepts related to explicit content and not all seven inappropriate concepts, thus having a smaller impact on specificity and fidelity.
@@
In contrast, ES has a negligible effect on specificity and fidelity, achieving CLPI and FID scores of 30.22 and 26.12, respectively, compared to the original SD-v1.4’s CLPI and FID scores of 31.31 and 25.25, indicating a compromise of about 3.7%. This impact is attributed to the default setting of a higher sanitization intensity ($\alpha=3.0$) in ES to enhance defensive performance, which moderately sacrifices generation quality. However, considering that ES improves effectiveness and robustness by approximately 25% to 40% over the best baseline, this slight quality trade-off is deemed acceptable. Figure 5 presents a visual comparison of clean generation content between ES and baseline methods, demonstrating that ES effectively preserves the original style and overall layout of images.
@@
Table IV: Generation quality of ES on the COCO-2017 dataset compared to the baseline methods.
Safeguard CLIP Score $\uparrow$ FID Score $\downarrow$
@@
Ours 30.22 26.12
5.4 Exploration on Hyperparameters
Figure 6: Performance of ES across different hyperparameters $\alpha$ (where $\beta=0.9$) and $\beta$ (where $\alpha=3.0$).
@@ 5.4 Exploration on Hyperparameters
The performance of ES is influenced by two key hyperparameters, $\alpha$ and $\beta$. The hyperparameter $\alpha$ is applied during the inference stage to control the sanitization strength of the prompt embedding, while $\beta$ is used during the training stage to handle potentially inappropriate synonyms in clean prompts $P_{c}$ within training sample pairs. We explore the impact of $\alpha$ and $\beta$ on ES by adjusting their values. Considering the computational cost of extensive parameter comparisons, we choose to measure the ER on the adversarial prompt dataset SP to evaluate defense performance, and use the COCO-2017 dataset to measure CLIP and FID scores for assessing generation quality.
As shown in Figure 6 we set the evaluation range of $\alpha$ and $\beta$ to $[0,4.5]$ and $[0,0.9]$ respectively, observing how ES’s performance changes as these parameters vary. The results indicate that as $\alpha$ decreases and $\beta$ increases, the ER generally rises, while CLIP and FID scores tend to worsen. This suggests that smaller values of $\alpha$ significantly enhance the sanitization effect of the prompt embedding, and larger values of $\beta$ enable better handling of target concept synonyms during training, guiding ES to learn through cleaner supervision signals. This leads to better defense performance but may come at the cost of generation quality. Our final experiments indicate that choosing $\alpha=3.0$ and $\beta=0.9$ provides an optimal balance between defensive performance and generation quality.
Moreover, $\alpha$ can be adjusted as a system variable post-deployment to further balance defense performance and generation quality based on practical needs. In Figure 7, we illustrate the generation results under different $\alpha$ values to verify the controllability of ES. Here, $\alpha=0$ represents the output of the original model. The first and second rows of the figure show sensitive content, where decreasing $\alpha$ results in progressively more appropriate clothing on figures, indicating that $\alpha$ effectively controls the defense strength. The third and fourth rows show benign content, where further decreasing $\alpha$ causes slight layout changes, which might explain the minor drop in generation quality.
6 Discussion
In this section, we analyze the interpretability and composability of ES, and compare the computational overhead introduced by ES. Last, we discuss some future works.
Figure 8: Examples illustrating the interpretability of ES.
(a) Interpretability Analysis. During the inference process of ES, S-Net provides interpretability by assigning scores between 0 and 1 to each token in the prompt. These scores measure the potential harmfulness, identifying the sources of inappropriate content generation. As shown in Figure 8, we analyze the interpretability of ES through examples of prompts and highlight three key observations.
Firstly, ES effectively erases target concepts. As illustrated in the first row of Figure 8, ES assigns high scores of 0.99 and 0.61 to the tokens ”nude” and ”person,” respectively. This is because ES suppresses ”nude” as a target concept during training to prevent explicit content, with the high score of 0.99 indicating that ES successfully identifies ”nude” and applies strong sanitization measures to erase it. Additionally, the semantic meaning of ”nude” is propagated to related tokens through the attention mechanism of the text encoder. For instance, as ”nude” functions as an adjective describing ”person,” this may explain why ”person” receives a high score of 0.61 in the context of ”nude.” Secondly, ES effectively recognizes synonyms of target concepts. As shown in the second and third rows of Figure 8, although ”topless” and ”bare” are not explicitly set as target concepts during training, they are synonyms of ”nude,” and ES assigns them high scores of 0.85 and 0.40, respectively. This indicates that ES demonstrates good generalization to synonyms of target concepts, effectively sanitizing them even if they are not explicitly defined during training. Thirdly, ES does not affect non-target concepts. As shown in the last two rows of Figure 8, ES tends to assign scores close to 0 to benign concepts. For example, ”A fashionable person,” which is unrelated to explicit content, receives low scores for all tokens. Notably, while the same token ”person” receives a high score in the context of ”nude”, it receives a low score in the context of ”fashionable,” demonstrating that ES does not assign identical scores to the same token indiscriminately but rather adjusts scores contextually. Additionally, the scores close to 0 indicate that the sanitization strength of ES is dynamically adjusted to the minimum, which explains why in the visualized results, ES’s benign outputs closely resemble those of SD-v1.4.
Table V: Performance of ES when combined with different T2I models and baseline methods.
Safeguard ER (%) $\uparrow$ CLIP $\uparrow$ FID $\downarrow$
I2P (sexual) SP RAB MMA COCO-2017
Ours + SD-v1.4 93.03 91.22 98.92 94.99 30.22 26.12
Ours + SD-v1.1 94.95 92.59 99.38 95.97 29.71 29.45
Ours + SD-v1.2 93.02 94.66 99.24 95.05 29.85 29.03
Ours + SD-v1.3 92.87 90.87 98.74 94.59 30.13 26.31
Ours + SD-v1.5 91.39 92.59 98.58 94.64 31.44 24.71
Ours + NP 93.62 92.08 98.90 94.82 29.34 36.33
Ours + SLD 94.06 93.45 99.27 95.74 30.15 32.46
Ours + ESD 96.14 97.24 99.13 97.81 26.89 36.23
Ours + CA 90.94 90.53 98.63 93.95 30.05 25.79
Ours + Safe-CLIP 93.94 92.25 99.02 95.10 29.63 26.44
Ours + SafeGen 93.72 93.11 98.97 95.45 30.41 25.72
(b) Compatibility Analysis. The compatibility of ES is reflected in two aspects. On one hand, ES is compatible with different T2I models because its training process only requires access to the text encoder and does not depend on other components of the T2I model. This allows ES to be easily transferred to various T2I models that share the same text encoder structure, such as transitioning from SD-v1.1 to SD-v1.5. On the other hand, ES is also compatible with existing protective measures, as it functions as an auxiliary module that is orthogonal to methods based on fine-tuning or adjustments to the generation process. This enables ES to be seamlessly integrated with other protective measures, thereby further enhancing the safety performance of T2I models. Table V presents the compatibility results of ES, where Ours+SD-v1.4 serves as the baseline performance of ES. The results indicate that ES is applicable to different versions of T2I models and can cooperate effectively with most protective measures, achieving improved defense performance. It is worth noting that performance may be lower than the baseline in certain configurations. For instance, the FID score of Ours+NP is suboptimal, and the ER score of Ours+CA is lower, which may be attributed to the limited effectiveness of the combined method itself, thereby impacting the overall performance.
@@ 6 Discussion
(c) Overhead Comparison. Considering that ES is a plug-and-play module, it may introduce additional computational overhead in the T2I generation process. To assess its computational burden, we compare the average time required to generate an image using ES and the baseline methods. Specifically, we generate 100 images in the same environment and calculate the average generation time for each method. As shown in Figure 9, ESD, CA, Safe-CLIP, and SafeGen are based on fine-tuned models, so they do not introduce additional computational overhead. In contrast, SLD[35], due to its complex noise-adding process, incurs a significantly higher computational cost, with an average time of approximately 4.52 seconds per image. In comparison, although ES is a plug-and-play module, its lightweight design (only 43.12M parameters in ES) ensures efficient inference, making the average generation time per image is nearly identical to the baseline model (SD-v1.4), without introducing extra computational burden. This demonstrates that ES maintains efficient computational performance while achieving SOTA-level defense capabilities, making it suitable for real-time applications in practical scenarios.
@@
(d) Limitations and Future Works. This study takes an important step toward achieving an ideal safety framework for T2I models. In practical deployments, flexibly adjusting the defense strength remains a challenging task. We design a residual skip-connection-based sanitization method that allows the defense strength to be adjusted freely by tuning the sanitization parameter $\alpha$. However, as a global hyperparameter, $\alpha$ uniformly controls the suppression effect across all target concepts, which may not meet the need for fine-grained control over different target concepts. Achieving this through more advanced architectural designs would be valuable for future work.
Moreover, we observe that ES essentially constructs a clean text embedding, as it operates as a vision-agnostic defense framework at the output of the text encoder. This implies that ES could potentially be applicable beyond T2I tasks, benefiting any task that relies on textual modalities. For instance, text-to-video generation tasks that use text embeddings to guide video creation [38], and text-to-image retrieval tasks that leverage text embeddings for image search [32], could all gain from ES’s protective capabilities. Future work could explore validating our defense technique across more task scenarios that utilize text embeddings.

### v2

@@ 6.1 RQ1: Defense Performance Evaluation
As shown in Table III, we first report the total number of inappropriate content generated by the base model (SD-v1.4 without any safeguards) in each dataset. For example, the original SD-v1.4 generates 91 instances of hateful content using the I2P dataset. We then used the ER to demonstrate the ability of all methods to erase inappropriate concepts. The results show that ES achieves SOTA level of effectiveness, with the best erasure performance for all categories in the I2P dataset. In terms of the overall effect of erasure in I2P, ES achieves the highest average ER of 74.77%, followed by NP with an average ER of 56.53%. This indicates that ES surpasses the current best baselines in effectiveness by approximately 20%. The visual comparison between ES and the baseline methods provided in Figure 4 further validates the effectiveness of ES.
@@
Additionally, ES also achieves SOTA robustness, showing the best average defense performance against three types of adversarial attacks. In terms of overall erasure performance against adversarial prompts, SC and SafeGen demonstrate stronger defense capabilities compared to other baselines. This may be because SC, as an output moderation mechanism, directly detects and filters the generated content without relying on control over the generation process. For SafeGen, its text-agnostic design makes it less affected by synonym transformations of target tokens. In contrast, ES achieves the highest average ER of 91.73%, indicating an excellent synonym generalization ability for target concepts and effective resistance to inappropriate hand-crafted or adversarial prompts. Figure 9 in Appendix C further demonstrates the defense effectiveness of the baseline method and ES for different exposed body parts, such as the female breast and male genitalia, and the results verify that ES effectively suppresses the generation of all six exposed body parts.
@@
Figure 4: [RQ1 & RQ2] Visual comparison of generated content between ES and eight safe generation baselines. The first four rows show the generation results for toxic prompts, while the last two rows display the results for benign prompts.
6.2 RQ2: Benign Generation Preservation
@@
We use CLIP and FID scores to evaluate the specificity and fidelity of ES, comparing it with the SD-v1.4 base model and eight safe generation baselines. Table IV reports the quantitative results of each method in terms of specificity and fidelity. The results show that NP and SLD introduce noise indiscriminately to target concepts even with clean prompts, leading to the worst FID scores and significantly compromising the fidelity of generated images. Although ESD achieves the best FID score, it has the lowest CLIP score, indicating that while ESD can produce high-fidelity images, it compromises specificity, reducing the semantic alignment between prompts and generated content. POSI and RR perform well in both specificity and fidelity, likely at the cost of effectiveness and robustness. SafeGen and Moderator suppresses only concepts related to explicit content and not all seven inappropriate concepts, thus having a smaller impact on specificity and fidelity.
@@
In contrast, ES has a negligible impact on specificity and fidelity, with CLIP and FID scores of 31.00 and 24.56, respectively, compared to the original SD-v1.4’s CLIP and FID scores of 31.31 and 25.15. This effect is due to ES’s default setting of a higher sanitization intensity ($\alpha$ = 3.0) to enhance defense performance, which slightly sacrifices the semantic consistency of benign text with images. Figure 4 provides a visual comparison of the clean generation content produced by ES and the baselines, showing that ES effectively preserves the original style and overall layout of the images.
@@
Table IV: [RQ2] Generation quality of ES on the COCO-2017 dataset compared to the safe generation baselines.
Method CLIP Score $\uparrow$ FID Score $\downarrow$
@@
RR 31.21 25.54
Moderator 31.57 25.38
Ours 31.00 24.56
6.3 RQ3: Interpretability Evaluation
Figure 5: [RQ3] Interpretability examples. Lines 1–3 show synonym substitutions, lines 4 and 6 demonstrate paraphrasing, and the last line presents an implicit expression.
During the inference process of ES, S-Net assigns a score to each token to measure its toxicity and identify the triggers for unsafe generations. As shown in  Figure 5, we demonstrate the interpretability of ES through a case study and explain how ES generalizes to inappropriate tokens that were not seen during training.
We highlight four key observations: (i) Firstly, ES effectively erases target tokens. As illustrated in the first row of Figure 5, ES assigns high scores of 0.99 and 0.62 to the tokens ”nude” and ”person,” respectively. This is because ES explicitly suppresses ”nude” as a target token during training to prevent explicit content, with a high score of 0.99 indicating that ES successfully identifies ”nude” and applies strong sanitization measures to erase it. Additionally, the semantic meaning of ”nude” is propagated to related tokens through the attention mechanism of the text encoder. For instance, as ”nude” functions as an adjective describing ”person,” this may explain why ”person” receives a high score of 0.62 in the context of ”nude.” (ii) Second, ES identifies inappropriate tokens that are not explicitly annotated during training. As shown in the second and third rows of  Figure 5, although “naked” and “topless” are not designated as target tokens during training, they are synonyms of “nude” and receive high scores of 0.90 and 0.44, respectively. This demonstrates that ES exhibits strong generalization and high coverage, effectively handling inappropriate synonyms. This supports our motivation for embedding-space sanitization: semantically similar tokens tend to be closer in the embedding space, facilitating ES ’s ability to generalize to unseen harmful expressions. (iii) Third, ES flexibly identifies inappropriate tokens based on context. As illustrated in the fourth row of  Figure 5, the word “suit” has different meanings depending on the context: in the phrase “birthday suit,” it implies nudity and receives a high score of 0.64, while in the context of “formal suit,” it refers to clothing and receives a low score of only 0.05. This likely benefits from the semantic encoding capability of the embedding representation. In the embedding space, “birthday suit” has a cosine similarity of 0.61 with “nude,” whereas “formal suit” has a similarity of only 0.13, indicating that ES can distinguish semantic differences based on context. (iv) Finally, ES detects inappropriate content expressed implicitly. As shown in the last row of  Figure 5, terms like “porn star,” while semantically distinct from explicit terms at the textual level, may still lead to similar visual outputs and thus exhibit certain relevance in the embedding space. ES captures this subtle connection and assigns a relatively high score of 0.30 to the token, demonstrating its ability to recognize implicit expressions of harmful content.
Overall, ES demonstrates strong generalization to synonyms and their implicit expressions, and does not assign fixed scores to the same token mechanically. Instead, it dynamically infers the true semantics based on context.
6.4 RQ4: Combinability Evaluation
Table V: [RQ4] Performance of ES when combined with different T2I models and baseline methods.
Method ER (%) $\uparrow$ CLIP $\uparrow$ FID $\downarrow$
SP RAB MMA COCO-2017
Ours + SD-v1.1 90.01 98.83 89.93 30.27 25.35
Ours + SD-v1.2 91.56 99.05 88.83 30.59 26.68
Ours + SD-v1.3 89.50 98.94 84.06 31.02 24.51
Ours + SD-v1.4 90.88 99.14 85.15 31.00 24.56
Ours + SD-v1.5 88.46 99.36 85.78 31.02 24.26
Ours + NP 93.11 98.52 93.15 29.84 30.86
Ours + SLD 90.87 98.49 90.79 30.27 29.27
Ours + ESD 95.00 99.44 98.56 28.85 25.88
Ours + POSI 92.25 98.72 89.75 30.08 25.33
Ours + Safe-CLIP 85.19 98.38 90.27 30.58 25.63
Ours + SafeGen 86.40 98.63 93.84 30.97 25.34
Ours + RR 89.15 98.83 81.87 30.97 24.91
Ours + Moderator 92.59 99.08 89.29 31.33 25.14
The compatibility of ES is reflected in two aspects. On one hand, ES is compatible with different T2I models because its training process only requires access to the text encoder and does not depend on other components of the T2I model. This allows ES to easily transfer to various T2I models using the same text encoder structure, such as from SD-v1.1 to SD-v1.5. On the other hand, ES is also compatible with existing safeguards because it functions as a plug-and-play model orthogonal to existing methods. This enables ES to seamlessly integrate with other safeguards, providing more comprehensive protection for T2I models. Table V shows the compatibility results of ES. When integrated with other safeguards, ES achieves significant performance improvements. For instance, in the Ours+ESD setting, ES achieves the best defense performance, while in the Ours+Moderator setting, it achieves the best specificity. This validates that ES is applicable to different T2I models and can effectively collaborate with other safeguards, further enhancing both defense performance and generation quality.
7 Discussion
Table VI: Ablation study results for ES.
Comparison Group ER (%) $\uparrow$ CLIP $\uparrow$ FID $\downarrow$
SP RAB MMA COCO-2017
A 79.69 99.02 57.30 31.23 25.06
B ($\lambda$ = 0.0) 93.46 98.80 96.14 27.07 29.10
B ($\lambda$ = 0.001) 85.88 98.99 90.05 30.67 25.05
B ($\lambda$ = 0.01) 90.88 99.14 85.15 31.00 24.56
B ($\lambda$ = 0.1) 79.86 98.97 47.98 31.25 24.42
C 78.66 64.21 48.36 31.23 25.08
(a) Ablation Studies. We design a series of ablation studies to evaluate the effectiveness and contribution of each component in ES. Detailed results are presented in Table VI.
  • Comparison Group A: Naive objective function. We first examine the performance of the naïve objective function. In this baseline, the optimization process excludes the purification of anchor prompts and uses the objective defined in Equation 6. Results show that this approach fails to generalize to a broader range of inappropriate synonyms. For instance, the interpretability score for the synonym “topless” is only 0.031, resulting in a significant drop in erasure rate. This supports our hypothesis in Equation 7, where simultaneously sanitizing the anchor prompt during optimization improves the generalization of ES to harmful tokens not explicitly annotated during training.
  • Comparison Group B: Effect of benign loss. Next, we assess the impact of the benign loss on ES performance. We vary the weight $\lambda=[0,0.001,0.01,0.1]$ in Equation 10 to control the strength of the benign constraint. When $\lambda=0$, the model achieves the strongest defensive performance but suffers a significant drop in generation quality. As $\lambda$ increases, generation quality improves while defensive effectiveness gradually decreases. These results guide us to select $\lambda=0.01$ as a balanced setting that maintains high generation quality with strong defense.
  • Comparison Group C: Semantically coherent synthetic data. Finally, we evaluate the contribution of semantic coherence in synthetic data. We employ the large language model Llama3-2 [13] to generate 50,000 semantically coherent toxic prompts. Specifically, for the seven target tokens defined in the training setup of Section 5, we prompt the LLM to compose natural sentences containing each token. We then generate corresponding anchor prompts by replacing the target tokens. Results indicate that semantic coherence in synthetic data does not lead to better performance, suggesting that ES is not sensitive to the coherence of training data.
Figure 6: Performance of ES across different $\alpha$.
(b) Controllability Analysis. In practical applications of ES, $\alpha$ serves as a tunable parameter after deployment, enabling dynamic trade-offs between defense performance and generation quality. For example, users with repeated violations can be assigned a larger $\alpha$ to further mitigate unsafe generations. As shown in Figure 6, we evaluate $\alpha$ in the range $[0,4.5]$. The results show that as $\alpha$ increases, ER tends to rise, while CLIP and FID scores slightly decrease, indicating that a larger $\alpha$ significantly enhances the purification of prompt embeddings. Figure 7 visualizes generation results under different $\alpha$ to demonstrate the controllability of ES. Here, $\alpha=0$ corresponds to the output of the original model. The first and second rows display sensitive content, where increasing $\alpha$ leads to more appropriate clothing, suggesting stronger defense. The third and fourth rows show benign content; increasing $\alpha$ causes slight layout changes, but the overall generation quality remains acceptable.
@@ 7 Discussion
(c) Overhead Comparison. To assess computational burden of ES, we compare the average time required to generate an image using ES and the baseline methods. Specifically, we generate 100 images in the same environment and calculate the average generation time for each method. As shown in Figure 8, fine-tuning-based approaches such as ESD do not introduce nearly as much additional computational overhead. POSI requires an additional language model to rewrite the prompts, increasing the computational costs by approximately 10%. SLD, due to its complex noise addition process, has the slowest generation time, averaging around 4.52 seconds per image. In comparison, although ES is a plug-and-play module, its lightweight design (only 43.12M parameters in ES) ensures efficient inference, making the average generation time per image nearly identical to the baseline model (SD-v1.4), without introducing extra computational burden. This shows that ES maintains efficient computational performance while achieving SOTA-level defense capabilities, making it suitable for application in real-time scenarios.
(d) Defense Internalization. Although ES, as a plug-and-play module, demonstrates strong safe generation capabilities in our threat model, its application in open-source scenarios may faces certain limitations. In open environments, adversaries may have the capability to bypass or disable ES, thereby escaping its defensive effectiveness. To address this challenge, we explore how to internalize the defense capabilities of ES into the model parameters. Inspired by the idea of knowledge distillation [12], we treat ES as a teacher model and distill its defensive ability into a text encoder $\mathcal{F}_{t,\omega}$ with parameters $\omega$. This process is formalized as follows:
$\mathcal{L}_{kd}=\min_{\omega}\left\|\mathcal{F}_{t,\omega}(P)-\hat{\text{Emb}}\right\|_{2}^{2},$ (11)
where $P$ denotes the input prompt, including both toxic and clean prompts, and $\hat{\text{Emb}}$ represents the sanitized embedding produced by ES (see Equation 7). Although some defensive capacity is inevitably lost during the distillation process, the overall performance still surpasses that of mainstream safety mechanisms. On three adversarial prompt datasets, the average erasure rate reaches 81.21%, while the CLIP and FID scores on the COCO dataset are 30.79 and 25.62, respectively. These results indicate that the defensive capabilities of ES can be effectively internalized into the model itself, providing security guarantees for safe generation in open-source settings.
@@
(e) Limitations and Future Works. This study takes an important step toward achieving an ideal safe generation framework for T2I models. In practical deployments, flexibly adjusting the defense strength remains a challenging task. We design a residual skip-connection-based sanitization method that allows the defense strength to be freely adjusted by tuning the sanitization parameter $\alpha$. However, $\alpha$ uniformly controls the suppression effect across all target concepts, which may not meet the need for fine-grained control over different target concepts. Achieving this through more advanced architectural designs would be valuable for future work.
Moreover, we observe that ES essentially constructs a clean text embedding, as it operates as a vision-agnostic defense framework at the output of the text encoder. This implies that ES could potentially be applicable beyond T2I tasks, benefiting any task that relies on textual modalities. For example, text-to-video generation tasks that use text embeddings to guide video creation [41], and text-to-image retrieval tasks that leverage text embeddings for image search [35], could all gain from ES’s protective capabilities. Future work could explore validating our defense technique in more task scenarios that utilize text embeddings.